OpenStudy (anonymous):

verify identity sinxsecxtanxcscx = sinx/1-sin2x

OpenStudy (gorv):

\[\sin x * \frac{ 1 }{ cosx }*\frac{ sinx }{cosx }*\frac{ 1 }{ sinx }=\frac{ sinx }{ \cos^{2}x }\]

OpenStudy (gorv):

cos2x=1-sin2x